Suppose that you buy a 3-month call for 100,000 widgets with a strike price of 7.26990 USD/Widget and pay a premium of 5 cents/Widget. Additionally, you enter into a short sale of 100,000 widgets at 6.64720 USD/Widget. At the time of maturity, the spot price is trading at 6.7137 USD/Widget. What are your profits/losses?

A) 21,990.00 USD

B) 15,340.00 USD

C) -11,650.00 USD

D) -18,360.00 USD
Answer #1

Payoff from buying a call option is maximum of (market price - strike price) or zero. If, at expiry, the market price is greater than the strike price then the buyer will exercise the call option and the payoff will be (market price - strike price).If, at expiry, the market price is lower than the strike price then the buyer will not exercise the call option and the payoff will be zero. So, the payoff at expiry will be either (market price - strike price) or zero, whichever is higher.

So, net profit (or loss) from buying the call option is cost of buying the premium + payoff

= premium + max(market price - strike price, 0)

Loss from buying the call option = number of widgets*(premium + max(spot price -strike price,0))

= 100,000*(0.05 + max(6.7137-7.2699, 0)) = 5000

Loss from short sale = number of widgets*(spot price - selling price)

= 100,000*(6.7137-6.64720) = 6,650

Total loss = -5,000 -6,500 = -11,500 (Option C)